When heating a 15 cm long square metal plate, its size increases by 0.05 cm.


Actions to be performed:


Answer: approximately how much did its area increase?


Determine how much its volume increases if it were a cubic plate.


If the plate is cooled and its side decreases by 0.05, how much did its area decrease?


If it is a cubic plate determined and it is cooled by decreasing its side by 0.05, determine how much its volume decreased?


1
Expert's answer
2022-03-02T16:51:46-0500

1. "\Delta x=0.05\ cm"


"A(x)=x^2"

"\Delta A=A(x_0+\Delta x)-A(x_0)=(x_0+\Delta x)^2-(x_0)^2"

"=(2x_0+\Delta x)\Delta x"

"\Delta A=(2(15cm)+0.05cm)(0.05cm)=1.5025{cm}^2"

The area increased by "1.5025\ {cm}^2."



"V(x)=x^3"

"\Delta V=V(x_0+\Delta x)-V(x_0)=(x_0+\Delta x)^3-(x_0)^3"

"=((x_0+\Delta x)^2+(x_0+\Delta x)x_0+(x_0)^2)\Delta x"

"\Delta V=((15.05cm)^2+15.05cm(15cm)+(15cm)^2)(0.05cm)"

"=33.862625 cm^3"

The volume increased by "33.862625 cm^3."


2. "\Delta x=-0.05\ cm"


"A(x)=x^2"

"\Delta A=A(x_0+\Delta x)-A(x_0)=(x_0+\Delta x)^2-(x_0)^2"

"=(2x_0+\Delta x)\Delta x"

"\Delta A=(2(15cm)-0.05cm)(-0.05cm)=-1.4975{cm}^2"

The area decreased by "1.4975\ {cm}^2."



"V(x)=x^3"

"\Delta V=V(x_0+\Delta x)-V(x_0)=(x_0+\Delta x)^3-(x_0)^3"

"=((x_0+\Delta x)^2+(x_0+\Delta x)x_0+(x_0)^2)\Delta x"

"\Delta V=((14.95cm)^2+14.95cm(15cm)+(15cm)^2)(-0.05cm)"

"=-33.637625 cm^3"

The volume decreased by "33.637625 cm^3."